One day, Kramer entered Jerry’s house 6 times more than George. If the difference between them is 15, how many times did George enter Jerry’s house.
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

One day, Kramer entered Jerry’s house 6 times more than George. If the difference between them is 15, how many times did George enter Jerry’s house.
-----------
K = G + 6G = 7G times
K - G = 15
------
7G - G = 15
6G = 15
G = 2.5 times
===================
The author of the problem is in the wrong business.
He/she doesn't understand math and/or English.

Answer by ikleyn(52799)

.
One day, Kramer entered Jerry’s house 6 times more than George. If the difference between them is 15,
how many times did George enter Jerry’s house.
~~~~~~~~~~~~~~~~~~~~~~~~~~

If to understand your condition LITERALLY, then the solution and the answer is absurdist. It is 2.5,  as Alan showed you in his solution.


The reason is that the phrase  

    "Kramer entered Jerry’s house 6 times more than George"  

in English  means / (is EQUIVALENT to)

    "Kramer entered Jerry’s house 7 times as many as George" .



There is no doubt, your intention was to say 

    "Kramer entered Jerry’s house 6 times as many as George".


Then the solution is

    6x - x = 15,

    5x = 15  =====>   x = 3   (ANSWER)

Solved.

------------------

    It is VERY COMMON error for those for whom English is not native language.

    In English,  this form works in TOTALLY DIFFERENT way than in many other languages !


Learn English from my posts !

And double your benefits in this way ! !


    It is a joke, of course.